OpenStudy (anonymous):

Which of the following is the best estimate of the number of centimeters in 1 yard? (Note 1 yard = 3 feet, 1 foot = 12 inches, 1 inch = 2.54 centimeters) a. 90 b. 180 c. 36 d. 72

OpenStudy (chosenmatt):

what did you come up with so far?

OpenStudy (anonymous):

1 inch = 2.54 cm 12 inches = 12*2.54 = 30.48 cm 1 foot = 12 inches 1 foot = 30.48 cms 3 feet = 3*30.48 cms = 91.44cms 1 yard = 3feet 1 yard = 91.44 cms